john can load his truck with peaches in 20 hours. his friend peter takes thrice as long to load John's truck with peaches. how long will it take them two working together to load the truck with peaches?
Answer by jorel555(1290)

John loads 1/20 of his truck per hour. Peter takes 3X as long, or 1/60 trucks per hour. So:
1/20+1/60=Amount they can fill per hour
3/60+1/60=4/60=1/15
Since they fill 1/15 truck per hour, it will take them 15 hours together to fill the truck!!!!!!!!!!!